# Large Learning Rates Simultaneously Achieve Robustness to Spurious Correlations and Compressibility

### v1

@@ 4.1 Datasets
We utilize two synthetic datasets from the literature to investigate the effects of LR on robustness to SCs and compressibility. These are the parity dataset proposed in [57] and the moon-star dataset proposed in [23]. The advantage of utilizing synthetic datasets is the clear and simple definition of simple vs. complex features they enable, as well as total control they afford over bias-conflicting sample ratio ($\rho$). In the parity dataset, the core and spurious features are binary vectors of size $C$ and $S$, with their parity bits (i.e. whether the vectors include odd number of 1’s) corresponding to the true label $y$ vs. spurious label $b$. Setting $C>S$ leads the spurious feature to be simpler than the complex feature. The moon-star dataset is a binary classification task where the classifier is expected to distinguish moon shaped objects from star shaped objects, with the spurious feature being the quadrant of the image on which the object is located. For both datasets we set bias-conflicting sample ratios as $\rho^{\textrm{train}}=0.1$ and $\rho^{\textrm{test}}=0.5$.
@@
These are arguably the most commonly used dataset types in research on SCs [63, 53, 31, 56, 57]. The reason for their popularity is the attractive combination they provide in the form of having relatively realistic inputs with a decent control over dataset properties such as bias-conflicting sample ratios and the complexity of the core vs. spurious features. We present the bulk of our results on four such datasets: Colored MNIST, Corrupted CIFAR-10, MNIST-CIFAR, and Double MNIST (see Fig. 2 for examples from each). The former two were proposed by [53], and have been frequently used in the literature. MNIST-CIFAR dataset is the extension of the domino dataset proposed by [65] to all 10 classes of MNIST and CIFAR-10, while Double MNIST has been designed by the authors for this paper. Each dataset includes some combination and/or modification of the well-known image datasets MNIST [38] and CIFAR-10 [34]. The core and spurious features for these datasets can be specified as digit shape vs. digit color for Colored MNIST, object vs. corruption type for Corrupted CIFAR-10, left digit vs. (brighter) right digit for Double MNIST, and CIFAR-10 vs. MNIST targets for MNIST-CIFAR. We set $\rho^{\textrm{train}}=0.025$ for Colored MNIST and Double MNIST datasets, and $\rho^{\textrm{train}}=0.1$ for Corrupted CIFAR-10 and MNIST-CIFAR, given the higher baseline difficulty of the latter two.
@@
As we extend our inquiry from SC datasets to naturalistic classification tasks, we investigate the effects of LR on model behavior using CIFAR-10, CIFAR-100, and ImageNet-1k datasets [34, 12]. These datasets include $32\times 32$ images of 10 and 100 classes of objects or animals, respectively. We use the traditional train-test splits in our experiments, consisting of 50000 and 10000 samples. ImageNet-1k consists of 1,331,167 color images scaled to $256\times 256$ belonging to 1000 classes. We use the traditional train-validation split with 1,281,167 and 50000 samples each.
@@ 4.2 Architectures, Training, & Implementation
To showcase the diversity of contexts in which LR has the aforementioned effects, we utilize a variety of architectures. These include fully connected networks (FCN) with ReLU activation, convolutional neural networks (CNN) with similar architectures to VGG11 [67], ResNet18 and ResNet50 [26], and Wide ResNet-101-2 [80]. We further include a large-scale vision transformer model, namely the Swin Transformer by [45]. Unless otherwise stated, all models are trained with a constant LR until 100% accuracy with no explicit L1/L2 regularization. Batch sizes are set to 16 for CelebA and Waterbirds experiments, and 100 for the remaining experiments. Our suppl. material provides additional details for our experiments and implementation, while showing that using different convergence criteria (i.e. training models longer after convergence), an LR annealing scheme, or using the popular Adam optimizer [30] leads to qualitatively identical results. An important exception to this scheme is the training of the Swin Transformer, where we utilize an AdamW optimizer combined with an ImageNet-1K pretraining initialization. This is done to test the applicability of our conclusions to realistic deployment scenarios. Unless otherwise noted, the results presented are the average of experiments run with at least 3 different random seeds. Since high LRs can lead to divergence and/or drastic performance loss at the extremes, we utilize LRs up to the point of divergence or drastic performance loss ($>25\%$ in test accuracy) among any one of the seeds. Lastly, since our results involve the examination of a number of variables simultaneously, in figures $y$-axes are normalized for each variable to highlight the effects of LR, and value ranges for all variables are separately provided in the supp. material for readability.
@@ 5 Results
[…]
@@
Do the observed effects of LR generalize to realistic scenarios? To make sure our results are not limited only to (semi-)synthetic data, we investigate the effect of LRs under a more realistic, up-to-date machine learning pipeline. As described in Sec. 4.2, we train a Swin Transformer on the CelebA dataset, and present its effects on worst-group performance, compressibility, and representation properties in Fig. 7. The results show that high LRs co-achieve robustness and compressibility in this scenario as well, further supporting our motivating hypothesis. See suppl. material for similar results on Waterbirds dataset.
@@
How does LR behave when compared to and combined with other hyperparameters? We then move on to investigate how large LRs compare to other important hyperparameters in creating these effects. The comparison set includes batch size, momentum, L1 and L2 regularization of the parameters, an alternative loss function designed for better exploitation of rare difficult examples (focal loss; FL [42]), and an optimization procedure for finding well-generalizing (wide) minima in the training loss landscape (adaptive sharpness aware minimization; ASAM [35, 20]). We repeat our experiments by varying the each hyperparameter in question (In the case of FL and ASAM these correspond to their central hyperparameters respectively, $\gamma$ and $\rho$). Fig. 6 demonstrates our results using a ResNet18 model: High LRs consistently facilitate a combination of OOD generalization, network compressibility, and core feature utilization, either on par with or surpassing other interventions. We also observe that other interventions (e.g. L1/L2 regularization) can be combined with high LRs to achieve even better trade-offs. Importantly, in Fig. 9 we highlight how different methods create robustness through different dynamics in the loss and feature utilization landscape, as discussed further in Sec. 6. See suppl. material for details, further results under other models and datasets as well as more background on FL and SAM.
@@ 6 On the mechanism of large LRs
We now investigate the mechanisms through which large LRs create their effects, and provide theoretical evidence for our findings. Numerous works indicate that simple spurious features are learned earlier than more complex core features [54, 57], regardless of LR [77, Theorem 4.1]: this produces a phase in learning where bias-aligned (BA) samples are predicted correctly, while bias-conflicting (BC) samples are misclassified. We find that at this stage, another effect of large LRs, namely fast norm growth of model parameters and logits, become important [^3] While norm growth under large LRs is well-established by previous research [64, 49, 46], its effects are underexplored in robust generalization.. This is because with cross-entropy (CE) loss, up-scaled logits under high LRs lead to confident mispredictions of BC samples. Due to the nonlinearty of CE, the mini-batch loss (and thus the gradient) is then increasingly dominated by mispredicted BC samples, corresponding to an implicit reweighting of the dataset in favor of BC samples. We now formalize this statement.
@@
The proof is deferred to the suppl. material, which also presents another proposition, showing that this results in an increase in the gradient norms to the subnetworks that utilize spurious vs. core features. This implies that large LRs generate strong gradients that discourage reliance on spurious features, whereas small LRs lack this pressure, allowing models to retain spurious subnetworks while memorizing BC samples. Figs. 12 and 13 demonstrate the consistency of our empirical findings with the implications of Prop. 1. Experiments with ResNet18, Fig. 12 demonstrate that unbiased test accuracy very strongly correlates with the maximum loss ratio (i.e. BC loss/total minibatch loss) encountered during training across datasets, providing support for our proposed mechanism (see suppl. material for similar results with additional datasets and models). Fig. 13 presents a visualization of training dynamics under a large vs. small LR ($\eta=0.001$ vs $0.1$). We observe that early focus on spurious features are weaned off under large LRs, as bias-conflicting samples dominate the loss due to confident mispredictions, both in the semi-synthetic Colored MNIST dataset and the standard classification CIFAR-10 dataset.
@@
We can interpret the large, disruptive updates to spurious subnetworks as a recurring lottery ticket type scenario [21], where spurious subnetworks are effectively reset by large gradient updates whenever they rely too much on spurious feature and lead to confident mispredictions, similar to the “catapult mechanism”, proposed by [39].
@@ 7 Conclusion
Robustness to SCs and compressibility are crucial requirements for advancing safe, trustworthy, and resource-efficient deep learning. In the absence of strong theoretical guarantees, our controlled experiments demonstrate the efficacy of LR as a significant implicit bias for addressing these concerns simultaneously. We show that high LRs jointly obtain robustness to spurious correlations, network compressibility, and favorable representation properties.
@@
Our work is far from an end-to-end, convergent explanation involving optimization dynamics, parameter loss landscapes, and representations. Future work should study how multiple/hierarchical SCs interact with training dynamics, and design explicit training procedures for more robust, efficient models.
@@ Acknowledgments
S. Zafeiriou and part of the research was funded by the EPSRC Fellowship DEFORM (EP/S010203/1) and EPSRC Project GNOMON (EP/X011364/1). T. Birdal acknowledges support from the Engineering and Physical Sciences Research Council [grant EP/X011364/1]. T. Birdal was supported by a UKRI Future Leaders Fellowship [grant MR/Y018818/1] as well as a Royal Society Research Grant [RG/R1/241402].
@@ A Extended Related Work
While previous research showed a positive relationship between compressibility and generalization through theoretical and empirical findings [3, 71, 70, 4, 6], it is much less clear how well this applies to OOD generalization. Indeed, existing research provides at best an ambivalent picture regarding the simultaneous achieveability of generalization, robustness, and compressibility [74, 17, 15, 19]. Various studies have highlighted the impact of large LRs on generalization [40, 39, 50], model compressibility [4], and representation sparsity [2]; making it a prime candidate for further investigation regarding its ability to facilitate compressibility and robustness in tandem. [28] point out how LRs in early training prevents the iterates from being stuck in narrow valleys in the loss landscape, where the curvature in certain directions are high. [39, 50] point out the importance of large LRs in early training, where basin-jumping behavior leads to better generalizing and/or flatter minima [29]. While [40, 2, 81] focus on the effect of large LRs on feature learning, [60] demonstrate the crucial role of spurious / opposing signals in early training, and how progressive sharpening [9, 41] of the loss landscape in the directions that pertain to the representation of these features lead to the eventual downweighting of such nonrobust features. [60] further observe that this is due to discrete nature of practical steepest ascent methods (GD, SGD), as it is not observed in gradient flow regime, implicating learning rate as a prime candidate for modulating this behavior.
@@ B Extended Discussion of Our Contributions
Our findings call into question the assumptions regarding the inductive biases of “default” SGD. We find that LR selection can change the unbiased test set accuracy by up to $\sim$50(See Table S1). This has two major implications: (i) Any method that relies on assumptions regarding default behavior of neural networks (e.g. [53, 43]) should consider the fact that the said defaults can vary considerably based on training hyperparameters, including but not limited to LR. (ii) Any proposed intervention for improving robustness to SCs should consider utilizing large LRs as a strong baseline against the proposed method.
@@
We observe a strong interaction between overparametrization and LR in robustness to SCs. Our findings show that (Table S1) the range of test accuracies that can be obtained by tuning LRs increase as the models get more expressive. For example, in Colored MNIST dataset, while the difference between the highest vs. lowest performing models is $\sim$4for a fully connected network (FCN), this increases to $\sim$23for ResNet18 and $\sim$31for ResNet50. This implies that while overparametrization indeed seems to play an important role in robustness to SCs as suggested by some previous work [63] (cf. [56]), this needs to be considered in the context of central training hyperparameters, as they can modulate this vulnerability to a great degree.
@@
As noted above, [56] argue that max-margin classification inevitably leads to exploitation of spurious features, and LRs might protect against SCs by creating models closer to a uniform margin solution. To support their conjecture, they present evidence that shows average losses incurred from bias-aligned (BA) vs. bias-conflicting (BC) samples are closer in large LR models compared to small LR models. However, we find that their finding do not generalize across different data distributions. Computing avg. loss for BA samples / avg. loss for BC samples, we find that in both Colored MNIST and Double MNIST datasets, low LR models produce average losses that are closer in ratio ($33.9\times 10^{-}$ 3 vs. $33.4\times 10^{-}$ 3 in Colored MNIST and $39.5\times 10^{-}$ 3 vs. $67.4\times 10^{-}$ 6 in Double MNIST). This highlights the importance of testing such claims across diverse settings, and emphasizes the need for novel and systematic explanations for the effect of large LRs on robustness to SCs.

### v2

@@ 4.1 Datasets
We utilize two synthetic datasets from the literature to investigate the effects of LR on robustness to SCs and compressibility. These are the parity dataset proposed in [60] and the moon-star dataset proposed in [22]. The advantage of utilizing synthetic datasets is the clear definition of simple vs. complex features they enable, as well as total control they afford over bias-conflicting sample ratio ($\rho$). In the parity dataset, the core and spurious features are binary vectors of size $C$ and $S$, with their parity bits (i.e. whether the vectors include odd number of 1’s) corresponding to the true label $y$ vs. spurious label $b$. Setting $C>S$ leads the spurious feature to be simpler than the complex feature. The moon-star dataset is a binary classification task where the classifier is expected to distinguish moon shaped objects from star shaped objects, with the spurious feature being the quadrant of the image on which the object is located. For both datasets we set bias-conflicting sample ratios as $\rho^{\textrm{train}}=0.1$ and $\rho^{\textrm{test}}=0.5$.
@@
These are arguably the most commonly used dataset types in research on SCs [65, 56, 32, 59, 60]. The reason for their popularity is the attractive combination they provide in the form of having relatively realistic inputs with a decent control over dataset properties such as bias-conflicting sample ratios and the complexity of the core vs. spurious features. We present the bulk of our results on four such datasets: Colored MNIST, Corrupted CIFAR-10, MNIST-CIFAR, and Double MNIST (see Fig. 2 for examples from each). The former two were proposed by [56], and have been frequently used in the literature. MNIST-CIFAR dataset is the extension of the domino dataset proposed by [67] to all 10 classes of MNIST and CIFAR-10, while Double MNIST has been designed by the authors for this paper. Each dataset includes some combination and/or modification of the well-known image datasets MNIST [39] and CIFAR-10 [35]. The core and spurious features for these datasets can be specified as digit shape vs. digit color for Colored MNIST, object vs. corruption type for Corrupted CIFAR-10, left digit vs. (brighter) right digit for Double MNIST, and CIFAR-10 vs. MNIST targets for MNIST-CIFAR. We set $\rho^{\textrm{train}}=0.025$ for Colored MNIST and Double MNIST datasets, and $\rho^{\textrm{train}}=0.1$ for the remaining two, given the higher baseline difficulty thereof.
@@
As we extend our inquiry from SC datasets to naturalistic classification tasks, we investigate the effects of LR on model behavior using CIFAR-10, CIFAR-100, and ImageNet-1k datasets [35, 11]. The former two include $32\times 32$ images of 10 and 100 classes of objects or animals, respectively. We use the traditional train-test splits in our experiments, consisting of 50000 and 10000 samples. ImageNet-1k consists of 1,331,167 color images scaled to $256\times 256$ belonging to 1000 classes. We use the traditional train-validation split with 1,281,167 and 50000 samples each.
@@ 4.2 Architectures, Training, & Implementation
To investigate the diversity of contexts in which LR has the aforementioned effects, we utilize a variety of architectures. These include fully connected networks (FCN) with ReLU activation, convolutional neural networks (CNN; a slightly simplified variant of VGG11 [69]), ResNet18 and ResNet50 [26], and Wide ResNet-101-2 [84]. We further include a vision transformer model, namely the Swin Transformer by [47]. Unless otherwise stated, all models are trained with a constant LR until 100% accuracy with no explicit L1/L2 regularization. Batch sizes are set to 16 for CelebA, Waterbirds, and ImageNet experiments, and 100 for the remaining experiments. An exception to the training scheme above is the training of the Swin Transformer on CelebA and Waterbirds datasets, where we utilize an AdamW optimizer combined with an ImageNet-1K pretraining initialization, to investigate the effects of LRs under more realistic training schemes. Our suppl. material provides additional details for our experiments and implementation, while showing that using different convergence criteria, optimizers (Adam [31], PSGD [42]), or LR schedulers (Cosine annealing [48], WSD [27]) leads to qualitatively identical, and sometimes stronger results. Unless otherwise noted, the results presented are the average of experiments run with at least 3 different random seeds. Since high LRs can lead to divergence and/or drastic performance loss at the extremes, we utilize LRs up to the point of divergence or drastic performance loss ($25\%$ in test accuracy) among any one of the seeds. Lastly, since our results involve the examination of a number of variables simultaneously, in figures $y$-axes are normalized for each variable to highlight the effects of LR, and value ranges for all variables are separately provided in the supp. material for readability. See our public repository [^3] https://github.com/mbarsbey/spurious_comp for additional implementation details.
@@ 5 Results
How does LRs affect robustness to SC and compressibility? We first investigate the effects of LR on compressibility and robustness to SCs, using the parity dataset in Fig. 3, with a fully connected network (FCN) with 3 hidden layers of width 200. The results show a clear effect of LR on robustness to SC and compressibility: Robustness and compressibility increases as a function of LR. The same figure also shows how the attributions to core, spurious, and noise pixels/features evolve for low and high LR models. Moving on to semi-synthetic SC datasets. Fig. 4 shows that across datasets and architectures, larger LRs lead to increased robustness to SCs, strongly supporting our central hypothesis. This increase in robustness is also accompanied by increased network and activation compressibility. Recall that the notion of $\kappa\textrm{-Prunability}$ we employ here is computed under an unbiased test set, emphasizing the high LRs effect on achieving these aims jointly. See suppl. material for additional details and similar results with moon-star and Corrupted CIFAR-10 datasets.
@@
Do the observed effects of LR generalize to larger models? To make sure our results are not limited to small models and (semi-)synthetic data, we investigate the effect of LRs using a larger, transformer architecture and naturalistic data. As described in Sec. 4.2, we train a Swin Transformer on the CelebA dataset, and present its effects on worst-group performance, compressibility, and representation properties in Fig. 7. The results show that high LRs co-achieve robustness and compressibility in this scenario as well, further supporting our motivating hypothesis. See suppl. material for similar results on Waterbirds dataset.
@@
How does LR behave when compared to and combined with other hyperparameters? We then move on to investigate how large LRs compare to other fundamental hyperparameters in a machine learning pipeline. The comparison set includes batch size, momentum, L1 and L2 regularization of the parameters, an alternative loss function designed for better exploitation of rare difficult examples (focal loss; FL [44]), and an optimization procedure for finding well-generalizing (wide) minima in the training loss landscape (adaptive sharpness aware minimization; ASAM [36, 19]). We repeat our experiments by varying the each hyperparameter in question (In the case of FL and ASAM these correspond to their central hyperparameters respectively, $\gamma$ and $\rho$). Fig. 6 demonstrates our results using a ResNet18 model: High LRs consistently facilitate a combination of OOD generalization, network compressibility, and core feature utilization, either on par with or surpassing other interventions. We also observe that other interventions (e.g. L1/L2 regularization) can be combined with high LRs to achieve even better trade-offs. Importantly, in Fig. 9 we highlight how different methods create robustness through different dynamics in the loss and feature utilization landscape, as discussed further in suppl. material, where we also share further details and results with other models and datasets, as well as more background on FL and SAM.
@@ 6 On the mechanism of large LRs
While plausible mechanisms through which large LRs create compressibility have been proposed [24], those through which large LRs can create robustness against SCs are much less explored, which we now investigate. Previous work indicates that simple spurious features are learned earlier than more complex core features [57, 60], regardless of LR [82, Theorem 4.1]: this produces a phase in learning where bias-aligned (BA) samples are predicted correctly, while bias-conflicting (BC) samples are misclassified. We find that at this stage, another effect of large LRs, namely fast norm growth of model parameters and logits, become important [^4] While norm growth under large LRs is well-established by previous research [66, 52, 49], its effects are underexplored in robust generalization.. This is because with cross-entropy (CE) loss, up-scaled logits under high LRs lead to confident mispredictions of BC samples. Due to the nonlinearty of CE, the mini-batch loss (and thus the gradient) is then increasingly dominated by mispredicted BC samples, corresponding to an implicit reweighting of the dataset in favor of BC samples. We now formalize this statement.
@@
The proof is deferred to the suppl. material, which also presents another proposition, showing how this can result in an increase in the gradient norms to the subnetworks that utilize spurious vs. core features. These imply that large LRs generate strong gradients that discourage reliance on spurious features, whereas small LRs lack this pressure, allowing models to retain spurious subnetworks while memorizing BC samples. Through experiments with ResNet18, Fig. 12 demonstrate that unbiased test accuracy very strongly correlates with the maximum loss ratio (i.e. BC loss/total minibatch loss) encountered during training across datasets, supporting our proposed mechanism (see suppl. material for similar results with additional datasets and models). Fig. 13 visualizes training dynamics under a large vs. small LR ($\eta=0.001$ vs $0.1$): early focus on spurious features are weaned off under large LRs, as BC samples dominate the loss due to confident mispredictions, both in Colored MNIST dataset and CIFAR-10 datasets.
@@
We can interpret these large, disruptive updates to the network under the outsized influence of bias-conflicting examples as a recurring lottery ticket type scenario [20], where spurious subnetworks are effectively reset by large gradient updates whenever they rely too much on spurious feature and lead to confident mispredictions, similar to the “catapult mechanism”, proposed by [40].
@@ 7 Conclusion
Robustness to SCs and compressibility are crucial requirements for advancing safe, trustworthy, and resource-efficient deep learning. In the absence of strong theoretical guarantees, our investigation demonstrates the efficacy of LR as a significant inductive bias for addressing these concerns simultaneously. We show that high LRs jointly obtain robustness to spurious correlations, network compressibility, and favorable representation properties.
@@
Our work is far from an end-to-end, convergent explanation involving optimization dynamics, parameter loss landscapes, and representations. Future work should also study how multiple/hierarchical SCs interact with training dynamics, and design explicit training procedures for more robust, efficient models.
@@ Acknowledgments
M. Barsbey was supported by the EPSRC Project GNOMON (EP/X011364/1). L. Prieto was supported by the UKRI Centre for Doctoral Training in Safe and Trusted AI [EP/S0233356/1]. S. Zafeiriou and part of the research was funded by the EPSRC Fellowship DEFORM (EP/S010203/1) and EPSRC Project GNOMON (EP/X011364/1). T. Birdal acknowledges support from the Engineering and Physical Sciences Research Council [grant EP/X011364/1]. T. Birdal was supported by a UKRI Future Leaders Fellowship [grant MR/Y018818/1] as well as a Royal Society Research Grant [RG/R1/241402].
@@ A Extended Related Work
While previous research showed a positive relationship between compressibility and generalization through theoretical and empirical findings [2, 73, 72, 3, 5], it is much less clear how well this applies to OOD generalization. Indeed, existing research provides at best an ambivalent picture regarding the simultaneous achieveability of generalization, robustness, and compressibility [77, 16, 14, 18]. Various studies have highlighted the impact of large LRs on generalization [43, 40, 53], model compressibility [3], and representation sparsity [1]; making it a prime candidate for further investigation regarding its ability to facilitate compressibility and robustness in tandem. [29] point out how large LRs in early training prevent the iterates from being stuck in narrow valleys in the loss landscape, where the curvature in certain directions is high. [40, 53] point out the importance of large LRs in early training, where basin-jumping behavior leads to better generalizing and/or flatter minima [30]. While [43, 1, 85] focus on the effect of large LRs on feature learning, [63] demonstrate the crucial role of spurious / opposing signals in early training, and how progressive sharpening [8, 78] of the loss landscape in the directions that pertain to the representation of these features lead to the eventual down-weighting of such non-robust features. [63] further observe that this is due to discrete nature of practical steepest ascent methods (GD, SGD), as it is not observed in gradient flow regime, suggesting learning rate as a prime candidate for modulating this behavior.
@@ B Extended Discussion of Our Contributions
  • Confident mispredictions of bias-conflicting samples play an important role in conferring robustness to models trained under large LRs.
@@
Our findings call into question the assumptions regarding the inductive biases of “default” SGD. We find that LR selection can change the unbiased test set accuracy by up to $\sim$50% (See Table S1). This has two major implications: (i) Any method that relies on assumptions regarding default behavior of neural networks (e.g. [56, 45]) should consider the fact that the said defaults can vary considerably based on training hyperparameters, including but not limited to LR. (ii) Any proposed intervention for improving robustness to SCs should consider utilizing large LRs as a strong baseline against the proposed method.
@@
We observe a strong interaction between overparametrization and LR in robustness to SCs. Our findings show that (Table S1) the range of test accuracies that can be obtained by tuning LRs increase as the models get more expressive. For example, in Colored MNIST dataset, while the difference between the highest vs. lowest performing models is $\sim$4% for a fully connected network (FCN), this increases to $\sim$23% for ResNet18 and $\sim$31% for ResNet50. This implies that while overparametrization indeed seems to play an important role in robustness to SCs as suggested by some previous work [65] (cf. [59]), this needs to be considered in the context of central training hyperparameters, as they can modulate this vulnerability to a great degree.
@@
As noted in the main paper, interventions developed to mitigate the effect of spurious correlations constitute two groups based on whether they assume access to spurious feature labels/annotations. Those that assume this, exploit this information to improve worst group or unbiased test set performance [64, 28]. In the absence of group annotations, other methods rely on assumptions about the nature of the spurious features, data distribution, and the inductive biases of the learning algorithms [45, 59, 74]. We highlight that the proposed mechanism in this paper, where large LRs cause BC sample losses to dominate acts as an implicit re-weighting, which makes it akin to two-stage methods like Learning from Failure [56]. This further highlights the importance of establishing the effects of core hyperparameters on robustness to spurious correlations - not only for a deeper understanding of the inductive biases of gradient-based learning under overparametrization, but also as strong baselines to compare against newly developed methods.
As noted above, [59] argue that max-margin classification inevitably leads to exploitation of spurious features, and LRs might protect against SCs by creating models closer to a uniform margin solution. To support their conjecture, they present evidence that shows average losses incurred from bias-aligned (BA) vs. bias-conflicting (BC) samples are closer in large LR models compared to small LR models. However, we find that their findings do not generalize across different data distributions. Computing avg. loss for BA samples / avg. loss for BC samples, we find that in both Colored MNIST and Double MNIST datasets, low LR models produce average losses that are closer in ratio ($3.9\times 10^{-}3$ vs. $3.4\times 10^{-}3$ in Colored MNIST and $9.5\times 10^{-}3$ vs. $7.4\times 10^{-}6$ in Double MNIST). This highlights the importance of testing such claims across diverse settings, and emphasizes the need for novel and systematic explanations for the effect of large LRs on robustness to SCs.